# Forming the local starburst galaxy Haro 11 through hydrodynamical merger simulations

### v1

@@ Abstract
Haro 11 is the closest known Lyman continuum leaking galaxy and serves as an important laboratory for studying the escape of Lyman continuum radiation. The galaxy is a metal-poor, starburst galaxy believed to be undergoing a merger that might help facilitate the escape of radiation. In this study, we carry out a large suite of numerical simulations of a merger between two disc galaxies, to study possible origins of Haro 11 and understand under which conditions various features of the galaxy are formed. By varying galaxy parameters describing the orbital configurations, masses, and their inclination, we perform a total of $\sim 500$ simulations. We demonstrate that a two-disc galaxy merger is able to reproduce key, observed features of Haro 11, including its morphology, gas kinematics, star formation history, and stellar population ages and masses. We also find that small parameter variations have minimal impact on the orbits and resulting galaxy properties. In particular, we present a fiducial Haro 11 model that produces the single observed tidal tail, the presence of three stellar knots, and inner gas morphology and kinematics. By performing mock observations, we compare with the results of observational data and discuss possible origins for various features. Furthermore, we present newly gathered observational data that confirms the presence of a stellar tidal tail with similar length and direction as our simulations.
@@
pubyear: 2025pagerange: Forming the local starburst galaxy Haro 11 through hydrodynamical merger simulationsA
@@ 1 Introduction
Understanding the impact of in-situ and external mechanisms on the evolution of galaxies remains a central challenge in extragalactic astronomy. Galaxy interactions and mergers are significant events in the formation history of a galaxy, being drivers of starbursts and morphological transformations, with ramifications for massive star cluster formation and the escape of ionising radiation (e.g. Kostyuk & Ciardi, 2024). Thus, knowing the specific scenario of a galaxy merger enables a deeper insight into the aforementioned physical mechanisms at work within the galaxy.
Blue compact dwarfs/galaxies (BCD/BCG) are compact and metal-poor galaxies undergoing a period of intense starburst resulting in especially young stellar populations dominating the blue part of the spectrum of the galaxy. Most BCGs have irregular morphologies (e.g. Bergvall & Östlin, 2002; Micheva et al., 2013), which would indicate current or recent close interaction and/or merger with a companion (e.g. Östlin et al., 2001; Adamo et al., 2011). This can also be inferred from tidal tails or gas debris, as well as periods of starburst. BCGs exhibit an extreme environment of stellar and cluster formation, and due to their low metallicity, compact size, and high SFR have been suggested to be analogues of young starburst galaxies in the high-$z$ Universe (e.g. Östlin et al., 2001; Adamo et al., 2011; Sirressi et al., 2022; Gao et al., 2022); believed to contribute a significant amount of the ionisation budget during the epoch of reionisation.
@@
Figure 1: Left: Image of Haro 11 from Adamo et al. (2010), taken by HST with the waveband filters 220W, 435W, and 814W. Right: Mock observations of the simulated galaxy, applying the same HST filters. Details of how this mock image was produced is described in Section 4.2. Both images are of the same spatial size and resolution.
@@
In order to numerically model a specific (observed) galaxy merger, a set of observational constraints that can be linked to orbital parameters are required. In particular, tidal tail features and stellar properties can yield strong constraints, but still suffer from degeneracy between parameters. However, previous numerical work of galaxy interactions have managed to pinpoint the initial orbital parameters of the progenitor galaxies that reproduce the general morphology and inherent galaxy properties. For example, simulations of the Antennae galaxies have managed to reproduce the morphology of the galaxies and their tidal tails, as well as the star formation, among other properties (Karl et al., 2010; Privon et al., 2013; Renaud et al., 2015; Lahén et al., 2018). In Östlin et al. (2015, hereafter Ö15), the authors provided a qualitative comparison between the Haro 11 galaxy and the Antennae galaxies. The Antennae galaxy is undergoing a merger and the authors show it to have similar morphology and kinematics as Haro 11. Although the Antennae galaxies are more massive and have two tidal tails, the fact that they have comparable kinematics and morphology suggests their orbital parameters might be similar.
@@
However, high-resolution simulations of specific galaxy mergers are still very limited. Even more rare are simulations of BCG/BCD (Bekki, 2008; Pearson et al., 2018; Zhang et al., 2020; Leicester et al., 2024), which are useful resources to interpret observational data and deepen our understanding about the complex physical mechanisms going on in these extreme galaxy environments. A benefit of simulations, particularly for BCGs, is the access to several observation angles, which allows study of anisotropy in radiation leaking. However, finding initial conditions that show a broad agreement across different observed features is not trivial and requires an extensive study of the parameters controlling the galaxy orbits.
@@
In this paper, we model the merger of two disc galaxies to reproduce the general morphology and properties of the Haro 11 galaxy as observed today, and discuss a possible origin of its formation. We employ the the hydrodynamical and $N$-body code RAMSES (Teyssier, 2002) to perform a scheme of low-resolution simulation tests ($\sim 100$ pc) of the interaction, varying the galaxy and orbital parameters until the merging galaxies’ morphology resembles the morphology of Haro 11 today. We compare with observational studies of the gaseous and stellar components, e.g. star formation rate and masses, to constrain the initial conditions our model. The simulation that best represents Haro 11 observational properties is then run at the higher spatial resolution of 36 pc. We present the results of this fiducial simulation and compare the successes and shortcomings of our models and discuss its applications. Furthermore, we present newly acquired deep imaging data that show the presence of a stellar tidal feature in the direction of the HI gas tail which have been predicted by the fiducial model, corroborating the good match.
@@ 2 Observational constraints
In this section, we briefly outline the key components of Haro 11 that were used to constrain and confirm the initial properties and orbital parameters of the progenitor galaxies. All of these observational constraints were considered during the formulation of our simulations’ initial conditions, but not all of them used, as matching every property in detail is not feasible. The method of determining our initial conditions is explained in Section 3.2.
@@ 2.1 Morphological appearance
A distinct feature of Haro 11 is its three bright stellar knots A, B, and C (see left plot of Figure 1 for an image of the galaxy). The morphology of the knots indicate that both B and C have a compact stellar core, while knot A appears as more spread out. This asymmetry in the stellar distribution has lead to speculations that the progenitor galaxies are dissimilar (see Ö15). Furthermore, this figure highlights another feature: a string of gas and stars that goes between knot B and A, which then loops back up and to the side. This feature has been dubbed an ’ear’ and is reminiscent of a disc arm, which could be interpreted as knot B being the centre of a disc and knot A forming within that disc (see e.g. Ö15, Menacho et al., 2019).
@@ 2.2 Recent star formation history
Haro 11 is currently undergoing a strong starburst with a star formation rate (SFR) upwards of $\lesssim 30\,{\rm\,M_{\odot}\,yr^{-1}}$ (Hayes et al., 2007; Adamo et al., 2010; Madden et al., 2013; MacHattie et al., 2014; Gao et al., 2022). Recently, Sirressi et al. (2022, see also ) used spectra to analyse the spectral energy distribution of the three knots in Haro 11 and determined that the energy budget is best modelled by three distinct stellar age populations: $1-4$ Myr, $4-40$ Myr, and $40-100$ Myr. From their analysis they estimate the stellar mass in each knot for the different ages at the range $10^{6}-10^{8}{\rm\,M_{\odot}}$.
One of the most significant features of Haro 11 is the leakage of Lyman continuum (LyC) radiation, making it the closest known LyC leaker (Bergvall et al., 2006). The origin of this radiation has predominantly come from knot B and C (Prestwich et al., 2015; Gross et al., 2021; Östlin et al., 2021; Komarova et al., 2024; Danehkar et al., 2024), with Komarova et al. (2024) reporting knot B to have a higher LyC production but lower escape fraction. While radiative transfer is beyond the scope of this study (but will be adressed in Ejdetjärn et al. in prep.), the presence of LyC radiation indicates an absence of gas in this region, toward knot C. This depletion could be the result of several mechanisms: gas being stripped during the interaction between galaxies (Le Reste et al., 2024; Kostyuk & Ciardi, 2024), starburst causing an increase in stellar feedback effects (e.g. Bergvall et al., 2013), or it may be linked to the intrinsic characteristics of the progenitor galaxies (e.g. dust-poor). Based on this, some authors have proposed that Haro 11’s progenitors could have been a dwarf disc galaxy and a less gas-rich galaxy, such as a dwarf spheroidal galaxy (e.g. Adamo et al., 2010; Östlin et al., 2015; Menacho et al., 2021).
@@ 2.3 Gas fraction
The gas mass around the central part of Haro 11 has been estimated for the molecular $M_{\rm H_{2}}=2.5\times 10^{8}-3.8\times 10^{9}{\rm\,M_{\odot}}$(Cormier et al., 2014; Gao et al., 2022), neutral HI $M_{\rm HI}=0.5\times 10^{9}{\rm\,M_{\odot}}$ (MacHattie et al., 2014; Pardy et al., 2016; Le Reste et al., 2024), and ionised $M_{\rm H\alpha}=1.4\times 10^{9}{\rm\,M_{\odot}}$ (Menacho et al., 2019) gas phases. The stellar mass is around $M_{*}=1.6\times 10^{10}{\rm\,M_{\odot}}$ (Ö15), which yields a total gas fraction $f_{\rm g}\sim 14-27\,\%$; depending on the molecular mass. The neutral HI mass was recently reassessed by Le Reste et al. (2024), who showed that previous observations had not been sensitive enough to detect the tidal tail, which they note makes up $45^{+19}_{-21}$ of the total HI mass, $M_{\rm HI}=1.11\times 10^{9}{\rm\,M_{\odot}}$. The detection of a stellar tidal tail could possibly change the galaxy’s gas fraction further. While the mass of the various gas phases provide weak constraints, due to the sensitivity to specific feedback and star formation models, the total gas mass and gas fraction are viable for us to match.
@@
Halo Mass $[{\rm 10^{9}\ M_{\odot}}]$ 200 167
@@
Gas disc Mass $[{\rm 10^{9}\ M_{\odot}}]$ 5.5 1.2
@@
Stellar disc Mass $[{\rm 10^{9}\ M_{\odot}}]$ 4.57 2.57
@@
Stellar bulge Mass $[{\rm 10^{9}\ M_{\odot}}]$ 4.0 2.28
@@
Position [kpc] Velocities [${\,\rm{km\,s^{-1}}}$ ] inclination
@@ 3 Numerical simulations and methods
In this section we outline the physical recipes for star formation and feedback, and describe our method for determining the initial properties and orbital parameters of the progenitor galaxies. The simulations presented in this work were performed with the hydrodynamics and $N$-body code RAMSES (Teyssier, 2002), which uses an Adaptive Mesh Refinement (AMR) grid method with an HLLC Riemann solver (Toro et al., 1994). The code solves the conservative Euler equations for the dynamics of the gaseous component using a 2nd order Godunov scheme. We adopt an ideal mono-atomic gas with adiabatic index $\gamma=5/3$. The Euler equations consider the hydrodynamics and the gravitational effects from stars, dark matter, and fluid self-gravity. Dark matter and stars only interact through gravity, as collisionless particles. We employ a sub-grid recipe for the formation of stars and their feedback, which we will describe briefly in the coming section and refer to Agertz et al. (2013, 2021) for details.
The aim of this study is to explore possible formation scenarios for the Haro 11 galaxy. To this end, we have performed roughly 500 low-resolution simulation tests of a merger between two disc galaxies, as part of an iterative trial-and-error parameter study to approach the orbital and galaxy parameters that agree with Haro 11 observations. To reduce the run time, no stellar feedback was included in the initial test simulations. For each simulation we varied various parameters relevant for the progenitor galaxy orbits (e.g. velocity) or their intrinsic properties (e.g. disc size) to make the simulation fit, by eye, with the morphology, SFR, kinematics, and stellar properties observed in Haro 11. Parameter changes that improved the model were kept and further iterated upon. However, there is no definitive best set of parameters as orbital and galactic parameters interact complexly, causing parametric degeneracies where significantly different sets of initial conditions can achieve similar results.
@@ 3.1 Star formation and feedback physics
Stellar feedback follows the subgrid recipe detailed in Agertz et al. (2013, see also ) and tracks the injection of momentum, energy, and metals of supernovae Type Ia, Type II, and stellar winds. Individual supernovae are considered resolved when the host cell is resolved by at least three cooling lengths, following Kim & Ostriker (2015), and is otherwise deposited in the neighbouring gas cells. Metallicity is traced through oxygen and iron, which are combined into a total metallicity according to the relative solar abundance (Asplund et al., 2009).
@@ 3.2.1 Initial conditions
In order to determine the initial mass budget, we assumed the galaxies have retained a mass baryon fraction $f_{\rm b}$ close to the cosmic baryon fraction $\Omega_{\rm b}$, i.e. $f_{\rm b}=(M_{\rm g}+M_{*})/(M_{\rm g}+M_{*}+M_{\rm DM})\approx 15.6\%$ (Planck Collaboration et al., 2020). We base the initial stellar mass on observational constraints and do abundance matching (following Figure 9 in Behroozi et al., 2019) to determine the mass of the dark matter halo. The total stellar mass of Haro 11 is on the order of a few $10^{10}{\rm\,M_{\odot}}$ (Östlin et al., 2001; Madden et al., 2013, ; Ö15). The resulting dark matter matches with the total dynamical mass $\sim 10^{11}{\rm\,M_{\odot}}$ of Haro 11, estimated by Ö15 through dynamical reasoning of velocity field observations in the outer part of the galaxy.
The initial gas mass of the galaxies was calculated as the total observed mass between the molecular, neutral, and ionised gas phase (see Section 2 for values). We then added $1.25\times 10^{9}{\rm\,M_{\odot}}$ to this mass, in order to offset the gas consumed through star formation during the simulation run (determined during test runs). Similarly, we removed the same amount of mass from the observed stellar mass to determine our initial stellar mass. Additionally, some of this gas mass budget is added to the circumgalactic medium (see below). The gas mass in the discs is added in the simulation as neutral gas, and any additional gas phases naturally form as the galaxy evolves. The initial metallicity of the gas discs was directly taken from observations ($7.912+{\rm[O/H]}=$ 7.9, Ö15) and translated to a total metallicity $Z=0.2\,Z_{\odot}$.
The simulation was started at a low spatial resolution to allow the initial conditions to relax and then increased after a few 100 Myr, with plenty of time before the first interaction. The maximum spatial resolution of the fiducial simulation is $\Delta x=36$ pc, and is within a 3003 kpc3 volume. The circumgalactic medium spans the entire simulation volume with an initial temperature $T\sim 10^{8}$ K and hydrogen number density $n_{\rm H}=10^{-4}\,{\rm cm}^{-3}$, which is within the densities observed of local galaxies and in simulations (Machado et al., 2018). A proper gas halo is necessary to produce realistic outflows and tidal tails, and to ensure the gas mass of the system matches the cosmic baryon fraction for the given stellar mass. However, the circumgalactic gas content varies widely between systems and is not well constrained. We constructed a tests with higher and lower circumgalactic densities, which is briefly discussed in Section 5.4.

### v2

@@ Abstract
Haro 11 is a metal-poor, starburst galaxy believed to be the result of an ongoing merger, which is shaping the properties of the galaxy. In this study, we carry out a large suite of numerical simulations of a merger between two disc galaxies, to study possible origins of Haro 11 and understand under which conditions various features of the galaxy are formed. By varying galaxy parameters describing the orbital configurations, masses, and their inclination, we perform a total of $\sim 500$ simulations. We demonstrate that a two-disc galaxy merger reproduces key, observed features of Haro 11, including its morphology, gas kinematics, star formation history, and stellar population ages and masses. In particular, we present a fiducial Haro 11 model that produces the single observed tidal tail, three stellar knots, and inner gas morphology and kinematics. The resulting orbit and galactic morphology are robust against small variations of the initial parameters. By performing mock observations, we compare with the results of observational data and discuss possible origins for various features. Furthermore, we present newly gathered observational data that confirms the presence of a stellar tidal tail with similar length and morphology as our simulations.
@@
pubyear: 2025pagerange: Forming the local starburst galaxy Haro 11 through hydrodynamical merger simulationsB
@@ 1 Introduction
Understanding the interplay of in-situ and external mechanisms on the evolution of galaxies remains a central challenge in extragalactic astronomy. Galaxy interactions and mergers are significant events in the formation history of a galaxy, found to be drivers of starbursts and large-scale morphological transformations (e.g. Mihos & Hernquist, 1996; Li et al., 2008; Lambas et al., 2012; Renaud et al., 2022), with possible ramifications for properties such as: massive star cluster formation (e.g. Kruijssen et al., 2012; Adamo et al., 2020), AGN fueling (Sharma et al., 2024; La Marca et al., 2024), and the escape of ionising radiation (e.g. Kostyuk & Ciardi 2024; Zhu et al. 2024; but see Mascia et al. 2025 for counter-examples). Thus, knowing the specific scenario of a galaxy merger enables a deeper insight into the observed properties and physical mechanisms at work within interacting galaxies.
Blue compact dwarfs/galaxies (BCD/BCG) are compact and metal-poor galaxies undergoing a period of intense starburst resulting in especially young stellar populations dominating the blue part of the spectrum of the galaxy. Most BCGs have irregular morphologies (e.g. Bergvall & Östlin, 2002; Micheva et al., 2013), which would indicate current or recent close interaction and/or merger with a companion (e.g. Östlin et al., 2001; Adamo et al., 2011). This can also be inferred from tidal tails or gas debris, as well as periods of starburst. BCGs exhibit an extreme environment of stellar and cluster formation, and due to their low metallicity, compact size, and high SFR have been suggested to be analogues of young starburst galaxies in the high-$z$ Universe (e.g. Östlin et al., 2001; Adamo et al., 2011; Sirressi et al., 2022; Gao et al., 2022).
@@
Figure 1: Left: Image of Haro 11 from Adamo et al. (2010), taken by HST with the waveband filters F220W, F435W, and F814W. Right: Mock observations of the simulated galaxy, applying the same HST filters. Details of how this mock image was produced is described in Section 4.2. Both images encompass a $\sim 8$ kpc box and have a resolution of $\sim 30-40$ pc.
@@
To model a specific (observed) galaxy merger, we require observational constraints that map to orbital parameters. In particular, tidal tail features and stellar properties are strong constraints, but with some degeneracy between parameters. However, previous numerical work of galaxy interactions have managed to pinpoint the initial orbital parameters of the progenitor galaxies that reproduce the general morphology and inherent galaxy properties. For example, simulations of the Antennae galaxies have managed to reproduce the morphology of the galaxies and their tidal tails, as well as the star formation, among other properties (Karl et al., 2010; Privon et al., 2013; Renaud et al., 2015; Lahén et al., 2018). In Östlin et al. (2015, hereafter Ö15), the authors provided a comprehensive comparison between the Haro 11 galaxy and the Antennae galaxies. The Antennae galaxies are undergoing a merger and the authors show it to have similar morphology and kinematics as Haro 11. Although the Antennae galaxies are more massive and have two tidal tails, the fact that they have comparable kinematics and morphology suggests their orbital parameters might be similar.
@@
However, high-resolution simulations of specific galaxy mergers are still very limited. Even more rare are simulations of BCG/BCD (Bekki, 2008; Pearson et al., 2018; Zhang et al., 2020; Leicester et al., 2024), which are useful resources to interpret observational data and deepen our understanding about the complex physical mechanisms going on in these extreme galaxy environments. However, finding initial conditions that show a broad agreement across different observed features is not trivial and requires an extensive study of the parameters controlling the galaxy orbits.
@@
In this paper, we model the merger of two disc galaxies to reproduce the general morphology and properties of the Haro 11 galaxy as observed today, and discuss a possible origin of its formation. We employ the hydrodynamical and $N$-body code RAMSES (Teyssier, 2002) to perform a suite of low-resolution simulation tests ($\sim 100$ pc) of the interaction, varying the galaxy and orbital parameters until the merging galaxies’ morphology resembles the morphology of Haro 11 today. We compare with observational studies of the gaseous and stellar components, e.g. star formation rate and masses, to constrain the initial conditions our model. The simulation that best represents Haro 11 observational properties is then run at the higher spatial resolution of 36 pc. We present the results of this fiducial simulation and compare the successes and shortcomings of our models and discuss its applications. Furthermore, we present newly acquired deep imaging data that show the presence of a stellar tidal feature in the direction of the HI gas tail which have been predicted by the fiducial model, corroborating the good match.
@@ 2 Observational constraints
In this section, we briefly outline the key components of Haro 11 that were used to constrain and confirm the initial properties and orbital parameters of the progenitor galaxies. All of these observational constraints were considered during the formulation of our simulations’ initial conditions, but not all of them were used, as matching every property in detail is not feasible. The method of determining our initial conditions is explained in Section 3.2.
@@ 2.1 Morphological appearance
A distinct feature of Haro 11 is its three bright stellar knots A, B, and C (see left plot of Figure 1 for an image of the galaxy). The morphology of the knots indicate that both B and C have a compact stellar core, while knot A appears as more spread out. This asymmetry in the stellar distribution has led to speculations that the progenitor galaxies are dissimilar (see Ö15). Furthermore, this figure highlights another feature: a string of gas and stars that goes between knot B and A, which then loops back up and to the side. This feature has been dubbed an ’ear’ and is reminiscent of a disc arm, which could be interpreted as knot B being the centre of a disc and knot A forming within that disc (see e.g. Ö15; Menacho et al., 2019).
@@ 2.2 Recent star formation history
Haro 11 is currently undergoing a strong starburst with a star formation rate (SFR) upwards of $\lesssim 30\,{\rm\,M_{\odot}\,yr^{-1}}$ (Hayes et al., 2007; Adamo et al., 2010; Madden et al., 2013; MacHattie et al., 2014; Gao et al., 2022), which are derived from different gas tracers and SED fitting methods but with an overall agreement. Recently, Sirressi et al. (2022, see also ) used spectra to analyse the spectral energy distribution of the three knots in Haro 11 and determined that the energy budget is best modelled by three distinct stellar age populations: 141-4 Myr, 4404-40 Myr, and 4010040-100 Myr. From their analysis they estimate the stellar mass in each knot for the different ages within the range $10^{6}-10^{8}{\rm\,M_{\odot}}$.
One of the striking features of Haro 11 is the leakage of Lyman continuum (LyC) radiation, making it the closest known LyC leaker (Bergvall et al., 2006). The origin of this radiation has predominantly come from knot B and C (Prestwich et al., 2015; Gross et al., 2021; Östlin et al., 2021; Komarova et al., 2024; Danehkar et al., 2024), with Komarova et al. (2024) reporting knot B to have a higher LyC emission and production but lower escape fraction. While radiative transfer is beyond the scope of this study (but will be addressed in Ejdetjärn et al. in prep.), the presence of LyC radiation indicates an absence of neutral gas and dust covering knot B and C. This depletion could be the result of several mechanisms: gas being stripped during the interaction between the progenitor galaxies (Le Reste et al., 2024), a displacement of stars relative to the gas during the close encounters (Ejdetjärn et al. in prep.), strong stellar feedback (driven by starburst) forming escape channels (e.g. Menacho et al., 2019), or it may be linked to the intrinsic characteristics of the progenitor galaxies (e.g. dust-poor). Based on the gas distribution around the knots, some authors have proposed that Haro 11’s progenitors could have been a dwarf disc galaxy and a less gas-rich galaxy, such as a dwarf spheroidal galaxy (e.g. Adamo et al., 2010; Östlin et al., 2015; Menacho et al., 2021).
@@ 2.3 Gas fraction
The gas mass around the central part of Haro 11 has been estimated for the molecular $M_{\rm H_{2}}=2.5\times 10^{8}-3.8\times 10^{9}{\rm\,M_{\odot}}$(Cormier et al., 2014; Gao et al., 2022), neutral HI $M_{\rm HI}=0.5\times 10^{9}{\rm\,M_{\odot}}$ (MacHattie et al., 2014; Pardy et al., 2016; Le Reste et al., 2024), and ionised $M_{\rm H\alpha}=1.4\times 10^{9}{\rm\,M_{\odot}}$ (Menacho et al., 2019) gas phases. The stellar mass is around $M_{*}=1.6\times 10^{10}{\rm\,M_{\odot}}$ (Ö15), which yields a total gas fraction $f_{\rm g}\sim 14-27\,\%$; depending on the molecular mass. The neutral HI mass was recently reassessed by Le Reste et al. (2024), who showed that previous observations had not been sensitive enough to detect the tidal tail, which they note makes up $45^{+19}_{-21}$ % of the total HI mass, $M_{\rm HI}=1.11\times 10^{9}{\rm\,M_{\odot}}$. The detection and mass-determination of a stellar tidal tail could possibly change the galaxy’s gas fraction further. The mass of specific gas phases provide weak constraints to our models, as the formation of e.g. molecular and ionised gas are highly sensitive to the specific feedback and star formation recipes. However, the total gas mass and gas fraction only depend on the star formation physics, which we confirm result in a SFR similar to observations (see Section 4.3).
@@
Halo Mass $\rm 10^{9}\ M_{\odot}$}] 200 167
@@
Gas disc Mass $\rm 10^{9}\ M_{\odot}$}] 5.5 1.2
@@
Stellar disc Mass $\rm 10^{9}\ M_{\odot}$}] 4.57 2.57
@@
Stellar bulge Mass $\rm 10^{9}\ M_{\odot}$}] 4.0 2.28
@@
Position [kpc] Velocities $\rm{km\,s^{-1}}$}] inclination
@@ 3 Numerical simulations and methods
In this section we outline the physical recipes for star formation and feedback, and describe our method for determining the initial properties and orbital parameters of the progenitor galaxies. The simulations presented in this work were performed with the hydrodynamics and $N$-body code RAMSES (Teyssier, 2002), which uses an Adaptive Mesh Refinement (AMR) grid method with an HLLC Riemann solver (Toro et al., 1994). The code solves the conservative Euler equations for the dynamics of the gaseous component using a second order Godunov scheme. We adopt an ideal mono-atomic gas with adiabatic index $\gamma=5/3$. The Euler equations consider the hydrodynamics and the gravitational effects from stars, dark matter, and fluid self-gravity. Dark matter and stars only interact through gravity, as collisionless particles. We employ a sub-grid recipe for the formation of stars and their feedback, which we will describe briefly in the coming section and refer to Agertz et al. (2013, 2021) for details.
The aim of this study is to explore possible formation scenarios for the Haro 11 galaxy. To this end, we have performed roughly 500 low-resolution simulation tests of a merger between two disc galaxies, as part of an iterative trial-and-error parameter study to approach the orbital and galaxy parameters that agree with Haro 11 observations. To reduce the run time, no stellar feedback was included in the initial test simulations. For each simulation we varied various parameters relevant for the progenitor galaxy orbits (e.g. velocity) or their intrinsic properties (e.g. disc size) to make the simulation provide a qualitative visual match to the observed morphology, SFR, kinematics, and stellar properties observed in Haro 11. Parameter changes that improved the model were kept and further iterated upon. However, there is no definitive best set of parameters as orbital and galactic parameters interact complexly, causing parametric degeneracies where significantly different sets of initial conditions can achieve similar results.
@@ 3.1 Star formation and feedback physics
Stellar feedback follows the subgrid recipe detailed in Agertz et al. (2013, see also ) and tracks the injection of momentum, energy, and metals of supernovae Type Ia, Type II, and stellar winds. Individual supernovae are considered resolved when the host cell is resolved by at least six cooling lengths, following Kim & Ostriker (2015), and is otherwise deposited in the neighbouring gas cells. Metallicity is traced through oxygen and iron, which are combined into a total metallicity according to the relative solar abundance (Asplund et al., 2009).
@@ 3.2.1 Initial conditions
In order to determine the initial mass budget, we assumed the galaxies have retained a mass baryon fraction $f_{\rm b}$ close to the cosmic baryon fraction $\Omega_{\rm b}$, i.e. $f_{\rm b}=(M_{\rm g}+M_{*})/(M_{\rm g}+M_{*}+M_{\rm DM})\approx 15.6\%$ (Planck Collaboration et al., 2020). We base the initial stellar mass on observational constraints and do abundance matching (following Figure 9 in Behroozi et al., 2019) to determine the mass of the dark matter halo. The total stellar mass of Haro 11 is on the order of a few $10^{10}{\rm\,M_{\odot}}$ (Östlin et al. 2001; Madden et al. 2013; Ö15). The resulting dark matter matches with the total dynamical mass $\sim 10^{11}{\rm\,M_{\odot}}$ of Haro 11, estimated by Ö15 through dynamical reasoning of velocity field observations in the outer part of the galaxy.
The initial gas mass of the galaxies was calculated as the total observed mass between the molecular, neutral, and ionised gas phase (see Section 2 for values). We then added $1.25\times 10^{9}{\rm\,M_{\odot}}$ to this mass, in order to offset the gas consumed through star formation during the simulation run (determined during test runs). Similarly, we removed the same amount of mass from the observed stellar mass to determine our initial stellar mass. Additionally, some of this gas mass budget is added to the circumgalactic medium (see below). The gas mass in the discs is added in the simulation as neutral gas, and any additional gas phases naturally form as the galaxy evolves. The initial metallicity of the gas discs was directly taken from observations ($12+{\rm[O/H]}=7.9$; Ö15) and translated to a total metallicity $Z=0.2\,Z_{\odot}$.
The simulation was started at a low spatial resolution to allow the initial conditions to relax and then increased after a few hundred Myr, with plenty of time before the first interaction. The maximum spatial resolution of the fiducial simulation is $\Delta x=36$ pc, and is within a 3003 kpc3 volume. The circumgalactic medium spans the entire simulation volume with an initial temperature $T\sim 10^{8}$ K and hydrogen number density $n_{\rm H}=10^{-4}\,{\rm cm}^{-3}$, which is within the densities observed of local galaxies and in simulations (Machado et al., 2018). A proper gas halo is necessary to produce realistic outflows and tidal tails, and to ensure the gas mass of the system matches the cosmic baryon fraction for the given stellar mass. However, the circumgalactic gas content varies widely between systems and is not well constrained. We constructed tests with higher and lower circumgalactic densities, which is briefly discussed in Section 5.4.